OpenStudy (anonymous):

se logarithmic differentiation to find dy/dx. y = (x^2 − 64/x^2 + 64) , x > 8 dy/ dx = can someone please help me?

myininaya (myininaya):

\[\ln(y)=\ln|\frac{x^2-64}{x^2+64}|\] \[\ln(y)=\ln|x^2-64|-\ln|x^2+64|\] \[\frac{y'}{y}=\frac{2x-0}{x^2-64}-\frac{2x+0}{x^2+64}\]

myininaya (myininaya):

\[y'=y(\frac{2x}{x^2-64}-\frac{2x}{x^2+64})\]

myininaya (myininaya):

\[y'=\frac{x^2-64}{x^2+64} (\frac{2x}{x^2-64}-\frac{2x}{x^2+64})\]

OpenStudy (anonymous):

log dif, is sweet

myininaya (myininaya):

\[y'=\frac{2x}{x^2+64}-\frac{2x(x^2-64)}{(x^2+64)^2}\]

myininaya (myininaya):

\[y'=\frac{2x(x^2+64)-2x(x^2-64)}{(x^2+64)^2}\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

wow your so good at this

OpenStudy (anonymous):

so thats the final answer?

myininaya (myininaya):

you might want to check my arithmetic

OpenStudy (anonymous):

so 126x/(x^2+64)^2 is the answer

myininaya (myininaya):

i don't have a calculator with me

OpenStudy (anonymous):

hmm 64*4 is 256

myininaya (myininaya):

yep it should be 256 lol

OpenStudy (anonymous):

omg lol

myininaya (myininaya):

\[y'=\frac{2(2(64)x)}{(x^2+64)^2}=\frac{4(64)x}{(x^2+64)^2}=\frac{256x}{(x^2+64)^2}\]

OpenStudy (anonymous):

hmm everything else seems ok..should is imply the bottom

OpenStudy (anonymous):

simpliify*

myininaya (myininaya):

what?

myininaya (myininaya):

simplify the bottom?

OpenStudy (anonymous):

could u quickly check the arimthmatic, using a online claculutor?

OpenStudy (anonymous):

like factor the bottom out?

myininaya (myininaya):

the bottom is factored

myininaya (myininaya):

do you multiply don't make something more ugly

myininaya (myininaya):

do you mean multiply*

OpenStudy (anonymous):

ok..so do u think thats right now..do u mind checking your work over i did but maybe you might catch a mistake

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

liek the bets way towrite it out..its a online quiz:(

myininaya (myininaya):

well i like the way we have our answer right now

myininaya (myininaya):

if you really want to check you work you can use wolfram

myininaya (myininaya):

i'm gonna go eat

OpenStudy (anonymous):

hmmm oki then.